# Hidden yet quantifiable: A lower bound for confounding strength using randomized trials

### v3

@@ 4.1 Datasets
We first benchmark tests and respective lower bounds with a synthetic distribution similar to [64, 33]. Here, the propensity score, the true unobserved confounding strength $\Gamma^{\star}$, and the correlation strength can be designed.
@@
We expand our benchmark using three real-world randomized trials: Hillstrom’s MineThatData Email data [27], the Tennessee STAR study [63] and the VOTE dataset [22]. In contrast to the synthetic experiments, these datasets involve real outcome functions, though the treatment assignment is still controlled.
We focus on Hillstrom’s dataset for clarity of presentation, and we refer the reader to Figure 5 for experiments on the other datasets showing similar trends. Hillstrom [27] focused on measuring the impact of an email campaign on the dollars spent by the recipients in the following two weeks. We first sample a small subset of the original trial, $D$, as our randomized trial, $D_{\operatorname{rct}}$. We can then subsample multiple observational studies from $D\setminus D_{\operatorname{rct}}$ sharing a fixed true confounding strength $\Gamma^{\star}$, i.e. $\mathbb{P}_{\operatorname{cnf}}$, but with a varying correlation between the hidden confounder $U$ and outcome $Y(1)$, i.e. $\mathbb{P}_{\operatorname{inv}}$.
Let us denote $X_{\mathrm{all}}$ as the vector of all observed covariates. While we cannot intervene on $\mathbb{P}_{\operatorname{inv}}(Y(1),Y(0)|X_{\mathrm{all}})$ as it is intrinsic to the dataset, we can generate multiple observational studies by partitioning $X_{\mathrm{all}}$ into unobserved $U$ and observed $X$ in different ways. For a given partitioning $X_{\mathrm{all}}=(U,X)$, the resulting $D_{\mathrm{os}}$ will have a specific $\mathbb{P}_{\operatorname{inv}}(Y(1),Y(0)|U)$ and hence correlation coefficient $\rho_{u,y}$. With each choice of $U$, we enforce a propensity score $\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=1\mid U)$ that satisfies $\mathcal{E}(\mathbb{P}^{\operatorname{os}},\Gamma^{\star})$ by subsampling $D\setminus D_{\operatorname{rct}}$. Finally, we remove $U$ to construct $D_{\mathrm{os}}$. Our subsampling approach is a variation of the methods presented in Keith et al. [37], Gentzel et al. [21] (see further details in Appendix B.2). Finally, we enforce 2.3 by excluding urban zip codes from the support of the randomized trial.
@@ 5 Real-world experiments
In essence, $\hat{\Gamma}_{\texttt{CT}}$ represents the minimum strength for which sensitivity analysis includes both positive and negative values of treatment effect, thereby invalidating the study results. Similar critical values have been proposed in the literature to assess the robustness of conclusions drawn from observational data, see e.g. VanderWeele and Ding [59], Jin et al. [33]. The most appropriate choice for the specific context should be determined by epidemiologists.
@@
For years, epidemiologists could not reach a consensus on the impact of hormone replacement therapy (HRT) on coronary heart disease and stroke based on the findings of the Women’s Health Initiative (WHI) study [1]. The WHI study included a randomized trial and an observational study that examined the impact of HRT on various cardiovascular events. While the observational study suggested that HRT had a protective effect against these outcomes, the randomized trial indicated the opposite. This discrepancy was recently resolved by identifying a strong unobserved confounder - the time $t$ since the start of HRT - and reanalyzing the data accordingly [58]. We now present evidence that our procedure can yield the same epidemiological conclusions and avoid issuing false alarms when the confounding is negligible.
@@ Results
In the case of stroke, both $\psi_{\texttt{sensi}}$ and $\psi_{\texttt{bin}}$ correctly flag the observational study, even when we adjust for the time since the start of treatment ($t=0$). This finding aligns with experts suggesting that additional unobserved confounding factors for stroke are still present after controlling for the time since the start of hormone replacement therapy [49].
@@ Acknowledgements
[…]
@@ A.1 Remaining proofs
Hence, if $\mathbb{P}_{H_{0}(\Gamma)}(\hat{T}_{\Gamma}^{-}\leq z_{\alpha/2})\leq\frac{\alpha}{2}+o_{\mathbb{P}}(1)$ and $\mathbb{P}_{H_{0}(\Gamma)}(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2})\leq\frac{\alpha}{2}+o_{\mathbb{P}}(1)$, the theorem follows from the union bound. For brevity, we only prove $\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}$ as the proof for $\hat{T}_{\Gamma}^{-}$ is analogous.
@@
$\sqrt{n_{\operatorname{rct}}}\left(\frac{1}{n_{\operatorname{rct}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)-\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right),\sigma^{2}+(\sigma^{+}_{\Gamma})^{2}-2\xi_{\Gamma}\right),\;\;\text{as}\;\;n_{\operatorname{rct}}\to\infty.$
@@
It remains to show that asymptotic normality also holds when we use the empirical estimate $\hat{\mu}_{\Gamma}^{+}$. To do so, we prove the following convergence in probability
$\left|\frac{1}{\sqrt{n_{\operatorname{rct}}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}_{\Gamma}^{+}(X_{i})\right|=o_{\mathbb{P}^{\operatorname{os}}}(1),\;\;\text{as}\;\;n_{\operatorname{rct}}\to\infty.$ (6)
Then, by Slutsky’s theorem, we have
@@
$\sqrt{n_{\operatorname{rct}}}\left(\hat{\mu}_{\Gamma}^{+}-\hat{\mu}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)-\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right),\sigma^{2}+(\sigma^{+}_{\Gamma})^{2}-2\xi_{\Gamma}\right),\;\;\text{as}\;\;n_{\operatorname{rct}}\to\infty\;\;\text{and}\;\;n_{\operatorname{os}}\to\infty.$ (7)
First, we observe that the mean of the LHS in Equation 6 converges to zero in probability:
$\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[\frac{1}{\sqrt{n_{\operatorname{rct}}}}\left|\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}_{\Gamma}^{+}(X_{i})\right|\right]\leq\sqrt{n_{\operatorname{rct}}}~{}\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[\left|\mu_{\Gamma}^{+}(X)-\hat{\mu}_{\Gamma}^{+}(X)\right|\right]\leq\sqrt{n_{\operatorname{rct}}}~{}\|\mu_{\Gamma}^{+}-\hat{\mu}_{\Gamma}^{+}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}=o_{\mathbb{P}^{\operatorname{os}}}(1),$
where in the last equality we have used $\|\mu_{\Gamma}^{+}-\hat{\mu}_{\Gamma}^{+}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}=O_{\mathbb{P}^{\operatorname{os}}}(n_{\operatorname{os}}^{-1/2})$ and $n_{\operatorname{rct}}\ll n_{\operatorname{os}}$. Next, we show that the variance term also converges to zero in probability, i.e.
$\operatorname{Var}_{\mathbb{P}^{\operatorname{rct}}}\left[\frac{1}{\sqrt{n_{\operatorname{rct}}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}_{\Gamma}^{+}(X_{i})\right]\leq\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[\left(\mu_{\Gamma}^{+}(X)-\hat{\mu}_{\Gamma}^{+}(X)\right)^{2}\right]=\|\mu_{\Gamma}^{+}-\hat{\mu}_{\Gamma}^{+}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}^{2}=o_{\mathbb{P}^{\operatorname{os}}}(1),$
and the statement in Equation 6 follows.
Therefore, by the consistency of $\hat{\sigma}^{2},(\hat{\sigma}^{+}_{\Gamma})^{2}$ and Slutsky’s theorem, we have
@@
where in the last line, we use Cauchy-Schwartz covariance inequality, i.e. $-\xi\leq|\xi_{\Gamma}|\leq\sigma^{+}_{\Gamma}\sigma$. Finally, by asymptotic normality established in Equation 7, we conclude that
$\lim_{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}\right)$ $\leq\lim_{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\frac{\sqrt{n_{\operatorname{rct}}}(\hat{\mu}^{+}_{\Gamma}-\hat{\mu})-\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)+\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right)}{\sqrt{(\sigma_{\Gamma}^{+})^{2}+\sigma^{2}-2\xi_{\Gamma}}}\leq z_{\alpha/2}\right)$
@@
Let $n=n_{\operatorname{rct}}+n_{\operatorname{os}}$ with fixed proportions, where $n_{\operatorname{rct}}=\rho n$ and $n_{\operatorname{os}}=(1-\rho)n$ for $\rho\in(0,1)$. Similarly to (1), by the central limit theorem and Lemma 3.2, it holds that
$\sqrt{n}\;\sum_{\begin{subarray}{c}(X_{i},T_{i},Y_{i})\in D_{\operatorname{rct}}\end{subarray}}\left(\frac{Y_{i}T_{i}}{\pi}-\frac{Y_{i}(1-T_{i})}{1-\pi}\right)w(X_{i})\xrightarrow{\mathcal{D}}\mathcal{N}\left(\mu\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}}\right),\sigma^{2}/\rho\right)\quad\text{as}\quad n\to\infty.$
@@
$\sqrt{n}\left(\begin{array}[]{c}\hat{\mu}_{\Gamma}^{+}\\ \hat{\mu}\end{array}\right)\stackrel{{\scriptstyle\mathcal{D}}}{{\rightarrow}}\mathcal{N}\left(\left(\begin{array}[]{l}\mu_{\Gamma}^{+}\\ \mu\end{array}\right),\left[\begin{array}[]{cc}(\sigma_{\Gamma}^{+})^{2}/(1-\rho)&0\\ 0&\sigma^{2}/\rho\end{array}\right]\right).$
@@
$\sqrt{n}\left(\hat{\mu}_{\Gamma}^{+}-\hat{\mu}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(\mu_{\Gamma}^{+}\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X}\right)-\mu\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right),\frac{(\sigma^{+}_{\Gamma})^{2}}{1-\rho}+\frac{\sigma^{2}}{\rho}\right)\quad\text{as}\quad n\to\infty.$
@@
$\frac{\mu^{+}_{\Gamma}-\hat{\mu}}{\sqrt{(\hat{\sigma}^{+}_{\Gamma})^{2}+\hat{\sigma}^{2}}}\xrightarrow{\mathcal{D}}\mathcal{N}\left(\mu_{\Gamma}^{+}\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X}\right)-\mu\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}}\right),1\right)\quad\text{as}\quad n\to\infty.$

### v4

@@ 4.1 Datasets
We first benchmark tests and respective lower bounds with a synthetic distribution similar to [yadlowsky2022bounds, jin2023sensitivity]. Here, the propensity score, the true unobserved confounding strength $\Gamma^{\star}$, and the correlation strength can be designed.
@@
We expand our benchmark using three real-world randomized trials: Hillstrom’s MineThatData Email data [hillstrom2008], the Tennessee STAR study [word1990state] and the VOTE dataset [gerber2008social]. In contrast to the synthetic experiments, these datasets involve real outcome functions, though the treatment assignment is still controlled.
We focus on Hillstrom’s dataset for clarity of presentation, and we refer the reader to Figure 5 for experiments on the other datasets showing similar trends. hillstrom2008 focused on measuring the impact of an email campaign on the dollars spent by the recipients in the following two weeks. We first sample a small subset of the original trial, $D$, as our randomized trial, $D_{\operatorname{rct}}$. We can then subsample multiple observational studies from $D\setminus D_{\operatorname{rct}}$ sharing a fixed true confounding strength $\Gamma^{\star}$, i.e. $\mathbb{P}_{\operatorname{cnf}}$, but with a varying correlation between the hidden confounder $U$ and outcome $Y(1)$, i.e. $\mathbb{P}_{\operatorname{inv}}$.
Let us denote $X_{\mathrm{all}}$ as the vector of all observed covariates. While we cannot intervene on $\mathbb{P}_{\operatorname{inv}}(Y(1),Y(0)|X_{\mathrm{all}})$ as it is intrinsic to the dataset, we can generate multiple observational studies by partitioning $X_{\mathrm{all}}$ into unobserved $U$ and observed $X$ in different ways. For a given partitioning $X_{\mathrm{all}}=(U,X)$, the resulting $D_{\mathrm{os}}$ will have a specific $\mathbb{P}_{\operatorname{inv}}(Y(1),Y(0)|U)$ and hence correlation coefficient $\rho_{u,y}$. With each choice of $U$, we enforce a propensity score $\mathbb{P}_{\operatorname{cnf}}^{\operatorname{os}}(T=1\mid U)$ that satisfies $\mathcal{E}(\mathbb{P}^{\operatorname{os}},\Gamma^{\star})$ by subsampling $D\setminus D_{\operatorname{rct}}$. Finally, we remove $U$ to construct $D_{\mathrm{os}}$. Our subsampling approach is a variation of the methods presented in keith2023rct, gentzel2021and (see further details in Appendix B.2). Finally, we enforce 2.3 by excluding urban zip codes from the support of the randomized trial.
@@ 5 Real-world experiments
In essence, $\hat{\Gamma}_{\texttt{CT}}$ represents the minimum strength for which sensitivity analysis includes both positive and negative values of treatment effect, thereby invalidating the study results. Similar critical values have been proposed in the literature to assess the robustness of conclusions drawn from observational data, see e.g. vanderweele2017sensitivity, jin2023sensitivity. The most appropriate choice for the specific context should be determined by epidemiologists.
@@
For years, epidemiologists could not reach a consensus on the impact of hormone replacement therapy (HRT) on coronary heart disease and stroke based on the findings of the Women’s Health Initiative (WHI) study [anderson2003implementation]. The WHI study included a randomized trial and an observational study that examined the impact of HRT on various cardiovascular events. While the observational study suggested that HRT had a protective effect against these outcomes, the randomized trial indicated the opposite. This discrepancy was recently resolved by identifying a strong unobserved confounder - the time $t$ since the start of HRT - and reanalyzing the data accordingly [vandenbroucke2009hrt]. We now present evidence that our procedure can yield the same epidemiological conclusions and avoid issuing false alarms when the confounding is negligible.
@@ Results
In the case of stroke, both $\psi_{\texttt{sensi}}$ and $\psi_{\texttt{bin}}$ correctly flag the observational study, even when we adjust for the time since the start of treatment ($t=0$). This finding aligns with experts suggesting that additional unobserved confounding factors for stroke are still present after controlling for the time since the start of hormone replacement therapy [prentice2005combined].
@@ Acknowledgements
The following appendices provide deferred proofs, experiment details, and ablation studies.
@@ A.1 Remaining proofs
Hence, if $\mathbb{P}_{H_{0}(\Gamma)}(\hat{T}_{\Gamma}^{-}\leq z_{\alpha/2})\leq\frac{\alpha}{2}+o(1)$ and $\mathbb{P}_{H_{0}(\Gamma)}(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2})\leq\frac{\alpha}{2}+o(1)$, the result follows from the union bound. For brevity, we only prove $\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}$ as the proof for $\hat{T}_{\Gamma}^{-}$ is analogous.
@@
$\sqrt{n_{\operatorname{rct}}}\left[\left(\frac{1}{n_{\operatorname{rct}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}\right)-\left(\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)-\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right)\right)\right]\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\sigma^{2}+(\sigma^{+}_{\Gamma})^{2}-2\xi_{\Gamma}\right),\;\;\text{as}\;\;n_{\operatorname{rct}}\to\infty.$
@@
It remains to show that asymptotic normality also holds when we replace the oracle average $\frac{1}{n_{\operatorname{rct}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})$ by the empirical estimator $\hat{\mu}_{\Gamma}^{+}$. To do so, define
$R_{n}:=\sqrt{n_{\operatorname{rct}}}\left(\frac{1}{n_{\operatorname{rct}}}\sum_{X_{i}\in D_{\operatorname{rct}}}\mu_{\Gamma}^{+}(X_{i})-\hat{\mu}_{\Gamma}^{+}(X_{i})\right).$ (6)
[…]
$\left|\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[R_{n}\mid D_{\mathrm{os}}\right]\right|\leq\sqrt{n_{\operatorname{rct}}}\,\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[|d_{n}(X)|\mid D_{\mathrm{os}}\right]\leq\sqrt{n_{\operatorname{rct}}}\,\|d_{n}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}=O_{\mathbb{P}^{\operatorname{os}}}\left(\sqrt{n_{\operatorname{rct}}/n_{\operatorname{os}}}\right)=o_{\mathbb{P}^{\operatorname{os}}}(1),$
where we used $\|d_{n}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}=O_{\mathbb{P}^{\operatorname{os}}}(n_{\operatorname{os}}^{-1/2})$ and $n_{\operatorname{rct}}/n_{\operatorname{os}}\to 0$. Similarly,
$\operatorname{Var}_{\mathbb{P}^{\operatorname{rct}}}\left[R_{n}\mid D_{\mathrm{os}}\right]\leq\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[d_{n}(X)^{2}\mid D_{\mathrm{os}}\right]=\|d_{n}\|_{L^{2}(\mathbb{P}^{\operatorname{rct}})}^{2}=O_{\mathbb{P}^{\operatorname{os}}}(n_{\operatorname{os}}^{-1})=o_{\mathbb{P}^{\operatorname{os}}}(1).$
Hence, by Chebyshev’s inequality, for every $\varepsilon>0$,
$\mathbb{P}^{\operatorname{rct}}\left(|R_{n}|>\varepsilon\mid D_{\mathrm{os}}\right)\leq\frac{1}{\varepsilon^{2}}\left(\operatorname{Var}_{\mathbb{P}^{\operatorname{rct}}}\left[R_{n}\mid D_{\mathrm{os}}\right]+\mathbb{E}_{\mathbb{P}^{\operatorname{rct}}}\left[R_{n}\mid D_{\mathrm{os}}\right]^{2}\right)=o_{\mathbb{P}^{\operatorname{os}}}(1).$
Since $0\leq\mathbb{P}^{\operatorname{rct}}\left(|R_{n}|>\varepsilon\mid D_{\mathrm{os}}\right)\leq 1$, this implies
$\mathbb{P}\left(|R_{n}|>\varepsilon\right)=\mathbb{E}_{\mathbb{P}^{\operatorname{os}}}\left[\mathbb{P}^{\operatorname{rct}}\left(|R_{n}|>\varepsilon\mid D_{\mathrm{os}}\right)\right]\longrightarrow 0,$
where $\mathbb{P}$ denotes the joint law of $(D_{\mathrm{os}},D_{\operatorname{rct}})$. Thus, $R_{n}=o_{\mathbb{P}}(1)$ under the joint law, and by the asymptotic normality established above and Slutsky’s theorem, we have, as $n_{\operatorname{rct}}\to\infty$ and $n_{\operatorname{os}}\to\infty$,
@@
$\sqrt{n_{\operatorname{rct}}}\left[\left(\hat{\mu}_{\Gamma}^{+}-\hat{\mu}\right)-\left(\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)-\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right)\right)\right]\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\sigma^{2}+(\sigma^{+}_{\Gamma})^{2}-2\xi_{\Gamma}\right).$ (7)
Finally, by the consistency of $\hat{\sigma}^{2},(\hat{\sigma}^{+}_{\Gamma})^{2}$ and Slutsky’s theorem, we have
@@
where in the last line, we use the Cauchy–Schwarz covariance inequality, i.e. $|\xi_{\Gamma}|\leq\sigma^{+}_{\Gamma}\sigma$. Further, under $H_{0}(\Gamma)$, it holds that $\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)-\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right)\geq 0$ by Lemma 3.1. Then, by asymptotic normality established in Equation 7, we conclude that
$\lim_{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\hat{T}_{\Gamma}^{+}\leq z_{\alpha/2}\right)$ $\leq\lim_{n_{\operatorname{rct}},n_{\operatorname{os}}\to\infty}\mathbb{P}_{H_{0}(\Gamma)}\left(\frac{\sqrt{n_{\operatorname{rct}}}\left((\hat{\mu}^{+}_{\Gamma}-\hat{\mu})-\mu_{\Gamma}^{+}\left(\mathbb{P}^{\operatorname{rct}}_{X}\right)+\mu\left(\mathbb{P}^{\operatorname{rct}}_{X},\mathbb{P}^{\operatorname{rct}}_{\operatorname{full}}\right)\right)}{\sqrt{(\sigma_{\Gamma}^{+})^{2}+\sigma^{2}-2\xi_{\Gamma}}}\leq z_{\alpha/2}\right)$
@@
Let $n=n_{\operatorname{rct}}+n_{\operatorname{os}}$ with fixed proportions, where $n_{\operatorname{rct}}=\rho n$ and $n_{\operatorname{os}}=(1-\rho)n$ for $\rho\in(0,1)$. By the central limit theorem and Lemma 3.2, it holds that
$\sqrt{n}\left(\frac{1}{n_{\operatorname{rct}}}\sum_{\begin{subarray}{c}(X_{i},T_{i},Y_{i})\in D_{\operatorname{rct}}\end{subarray}}\left(\frac{Y_{i}T_{i}}{\pi}-\frac{Y_{i}(1-T_{i})}{1-\pi}\right)w(X_{i})-\mu\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}}\right)\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\sigma^{2}/\rho\right)\quad\text{as}\quad n\to\infty.$
@@
$\sqrt{n}\left(\begin{array}[]{c}\hat{\mu}_{\Gamma}^{+}-\mu_{\Gamma}^{+}\\ \hat{\mu}-\mu\end{array}\right)\stackrel{{\scriptstyle\mathcal{D}}}{{\rightarrow}}\mathcal{N}\left(\left(\begin{array}[]{l}0\\ 0\end{array}\right),\left[\begin{array}[]{cc}(\sigma_{\Gamma}^{+})^{2}/(1-\rho)&0\\ 0&\sigma^{2}/\rho\end{array}\right]\right).$
@@
$\sqrt{n}\left(\hat{\mu}_{\Gamma}^{+}-\hat{\mu}-\mu_{\Gamma}^{+}\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X}\right)+\mu\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}}\right)\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\frac{(\sigma^{+}_{\Gamma})^{2}}{1-\rho}+\frac{\sigma^{2}}{\rho}\right)\quad\text{as}\quad n\to\infty.$
@@
$\frac{\hat{\mu}^{+}_{\Gamma}-\hat{\mu}-\mu_{\Gamma}^{+}\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X}\right)+\mu\left(\mathbb{P}^{\widetilde{\mathrm{os}}}_{X},\mathbb{P}^{\operatorname{os}}_{\operatorname{full}}\right)}{\sqrt{(\hat{\sigma}^{+}_{\Gamma})^{2}+\hat{\sigma}^{2}}}\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,1\right)\quad\text{as}\quad n\to\infty.$